OpenStudy (lgbasallote):

A Chinese restaurant charges a standard price for its "family dinner" consisting of any four different dishes on the menu. The restaurant's advertisement claims that "over 300 different family dinners are possible". If this is true, what must be the least number of dishes listed on the menu?

OpenStudy (anonymous):

any options available?

OpenStudy (unklerhaukus):

permutations or combinations?

OpenStudy (lgbasallote):

no idea

OpenStudy (lgbasallote):

but it is in that domain

OpenStudy (lgbasallote):

based on the question i would say combination

OpenStudy (anonymous):

you have the answer? so i can check

OpenStudy (unklerhaukus):

does a dinner have to include three items, ? can any of there be blank

OpenStudy (lgbasallote):

with all due respect @Omniscience if i give an answer i might get a random solution that gives the answer only by coincidence. So, i'd like to know the right solution so it might be best to keep the answer

OpenStudy (unklerhaukus):

i have mis-read the question

OpenStudy (unklerhaukus):

eleven ?

OpenStudy (lgbasallote):

how did you get 11?

OpenStudy (lgbasallote):

i know i have to do \[_n C _4 \ge 300\] but how do i solve for n?

OpenStudy (unklerhaukus):

well taking combinations \[C^k_4>300\] the smallest value of k that make the combination over 300 is k=11 \[C^{10}_4=210\] \[C^{11}_4=330\]

OpenStudy (unklerhaukus):

um i just tried some numbers , not exactly a great method

OpenStudy (unklerhaukus):

i dont think there is a inverse for combinations,

OpenStudy (lgbasallote):

hmm makes sense...

OpenStudy (lgbasallote):

so there's no other way to solve for n?

OpenStudy (unklerhaukus):

\[^nC_4\geq300\] \[\frac{n!}{4!(n-4)!}\geq300\] \[\frac{n(n-1)(n-2)(n-3)(n-4)!}{(n-4)!}\geq300\times4!\] \[{n(n-1)(n-2)(n-3)}\geq300\times4!\]